A triangle is cut from a rectangle. The height of a triangle is half of the unknown side length S. The area of the shaded region is 84 square inches. Write an equation you can use to find the side lengt S.
The long side of rectangle is 14. The short side is S. The base of the cut out triangle is 14.

Answer by addingup(3677)

area of the triangle:
b*h/2
area of your triangle:
14*s/2 = ? what is the "shaded area"? Is it the area of the triangle or something else?
IF 84 is the area of the triangle:
14*s/2 = 84
14*s = 168
s = 14
Now, you say that the side of the triangle is 1/2s, so if that's the case the side of the triangle would be 1/2s = 14/2 = 7